# Benchmarking Algorithms for Federated Domain Generalization

### v1

@@ Appendix D Supplementary Results
FedAvg-ERM $0.8389$ $0.8766$ $\mathit{0.9144}$ $\mathbf{0.8958}$ $\mathit{0.2727}$ $\mathbf{0.1762}$ $\mathbf{0.1241}$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $0.3134$ $0.0718$ $0.0924$
DeepCoral $\mathbf{0.9216}$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\mathbf{0.9395}$ $\mathbf{0.9437}$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8325$ $0.8407$ $0.8568$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8643$ $0.8787$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.8868}$ $0.8880$ $\mathit{0.8952}$ $0.1445$ $0.1023$ $0.0503$
FedADG - $0.0915$ $0.0732$ $0.0592$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1180$ $0.1341$ $0.1138$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8180$ $0.0271$ $0.0259$ $0.0081$
More results on CivilComments and Py150 dataset.
Here we report the DG accuracy on CivilComments and Py150 Table 6 using the in-domain validation. We could observe that most of the methods perform well on CivilComments and Py150. This could be attributed to the use of pretrained models. Additionally, the utilization of pretrained models may explain the relatively high value of $R_{\text{DG}}$, as NLP pretrained models have already been exposed to various domains.
Table 7: Test accuracy on CivilComments and Py150 datasets with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $0.6405$
DG Adapted IRM $\mathit{0.6213}$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $0.6038$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $0.5981$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $\mathit{0.6565}$
DeepCoral $0.5827$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $\mathbf{0.6383}$ $0.6028$ $0.4804$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - $0.3603$ $0.3389$ $0.3188$ $0.0915$ $0.0860$ $0.0587$
FedGMA - $0.3341$ $0.3157$ $0.3195$ $0.6628$ $0.3737$ $0.3576$
Results on FEMNIST dataset.
As mentioned in the main paper, we include the FEMNIST dataset here as reference. We could observe from Table 8 that $\lambda$ does not influence the final DG accuracy, and either in-domain validation or held-out-domain validation does not affect the final DG accuracy. This indicates the lacking of the statistical heterogeneity across different domains. we observe that changing of $\lambda$ does not significantly affect the convergence. Most of which does not converge to the centralized counterpart’s performance, this is due to the challenge coming from large number of clients where $R_{\text{FL}}=0.980<1$.
Table 8: Test accuracy on FEMNIST dataset with in-domain validation and held-out validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
FEMNIST (In-Domain) FEMNIST (Held-Out-Domain)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
B
FedAvg-ERM $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$ $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$
DG Adapted IRM $0.8442$ $0.8303$ $0.8389$ $0.8328$ $0.8442$ $0.8303$ $0.8389$ $0.8328$
Fish $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$ $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$
Mixup $0.8337$ $0.8275$ $0.8168$ $0.8161$ $0.8337$ $0.8275$ $0.8168$ $0.8161$
MMD $0.8442$ $0.8430$ $0.8323$ $0.8293$ $0.8442$ $0.8430$ $0.8323$ $0.8293$
DeepCoral $0.8457$ $0.8403$ $0.8328$ $0.8513$ $0.8457$ $0.8403$ $0.8328$ $0.8513$
GroupDRO $0.8413$ $0.8447$ $0.8139$ $0.8052$ $0.8413$ $0.8447$ $0.8139$ $0.8052$
FL FedProx - $0.8448$ $0.8440$ $0.8425$ $0.8448$ $0.8440$ $0.8425$
Scaffold - $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$ $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$
FDG FedDG - $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$ $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$
FedADG - $0.0549$ $0.0557$ $0.0548$ $0.0549$ $0.0557$ $0.0548$
FedSR - $0.8320$ $0.8316$ $0.8306$ $0.8320$ $0.8316$ $0.8306$
FedGMA - $\mathit{0.8486}$ $0.8395$ $0.8371$ $0.8486$ $0.8395$ $0.8371$
Appendix E Gap Table
@@
We list the gap table in Table 9 for summarizing the current DG algorithms performance gap w.r.t FedAvg-ERM in the FL context, in particular, positive means it outperforms FedAvg-ERM, negative means it is worse than FedAvg-ERM. It can be seen that in the on the simple dataset, the best DG migrated from centralized setting is better than FedAvg-ERM. In the domain separation case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting. However, they fail in harder datasets. In the hardest setting, currently the Federated methods dealing with data heterogeneity performs the best. It is worth noting that while federated learning methods that address client heterogeneity perform better than other methods, they still fall short of achieving centralized empirical risk minimization (ERM). This highlights the need for future research and development of DG methods in the FL regime.
Table 9: Gap Table: Current progress in solving Federated DG with respect to FedAvg-ERM.
Domain Separation
@@
Adapted centralized methods $+0.0629$ $-0.0981$ $+0.0421$ $-0.0062$
Federated methods $+0.0060$ $+0.0458$ $+0.0421$ $-0.0062$
FDG FedDG $+0.0468$ $-0.0775$
FedADG $-0.6874$ $-0.2171$
FedSR $-0.7520$ $-0.2164$ $-0.0326$ $-0.5912$
FedGMA $+0.0054$ $-0.1263$ $-0.0588$ $-0.0199$
Appendix F Training Time, Communication Rounds and local computation
@@
Table 10: Wall-clock Training time per communication (unit: s).
PACS IWildCam CivilComments Py150 FEMNIST
$C=100$ $C=243$ $C=100$ $C=100$ $C=100$
@@
FedAvg-ERM $143$ $6301$ $3958$ $6566$ $262$
DG Adapted IRM $147$ $6454$ $4085$ $7089$ $297$
Fish $148$ $7072$ $5483$ $7770$ $324$
Mixup $144$ $6294$ $264$
MMD $144$ $6663$ $4024$ $7603$ $287$
DeepCoral $144$ $6597$ $3901$ $7212$ $287$
GroupDRO $145$ $9311$ $4690$ $8121$ $307$
FL FedProx $169$ $6921$ $4502$ $6513$ $288$
Scaffold $167$ $6876$ $4421$ $6389$ $281$
FDG FedDG $352$ $32172$ $989$
FedADG $181$ $11094$ $907$
FedSR $151$ $7136$ $4403$ $8020$ $280$
FedGMA $143$ $6795$ $4525$ $6545$ $261$

### v2

@@ Appendix F Supplementary Results
FedAvg-ERM $\mathbf{0.854\pm 0.001}$ $0.837\pm 0.003$ $0.838\pm 0.002$ $0.834\pm 0.001$
DG Adapted IRM $0.844\pm 0.001$ $0.832\pm 0.007$ $0.822\pm 0.005$ $0.821\pm 0.004$
Fish $\mathit{0.849\pm 0.002}$ $0.833\pm 0.004$ $0.829\pm 0.003$ $0.826\pm 0.008$
Mixup $0.834\pm 0.002$ $0.828\pm 0.009$ $0.821\pm 0.008$ $0.816\pm 0.005$
MMD $0.844\pm 0.003$ $0.843\pm 0.006$ $0.832\pm 0.007$ $0.829\pm 0.006$
DeepCoral $0.846\pm 0.002$ $0.840\pm 0.005$ $0.836\pm 0.002$ $0.851\pm 0.003$
GroupDRO $0.841\pm 0.004$ $0.835\pm 0.005$ $0.814\pm 0.002$ $0.805\pm 0.003$
FL FedProx - $\mathbf{0.865\pm 0.001}$ $\mathbf{0.864\pm 0.001}$ $\mathbf{0.863\pm 0.001}$
Scaffold - $\mathit{0.860\pm 0.002}$ $\mathit{0.859\pm 0.002}$ $\mathit{0.858\pm 0.002}$
AFL - $0.629\pm 0.010$ $0.593\pm 0.037$ $0.584\pm 0.020$
FDG FedDG - $0.859\pm 0.002$ $0.857\pm 0.002$ $0.856\pm 0.002$
FedADG - $0.817\pm 0.026$ $0.800\pm 0.061$ $0.785\pm 0.025$
FedSR - $0.832\pm 0.015$ $0.832\pm 0.020$ $0.831\pm 0.014$
FedGMA - $0.849\pm 0.010$ $0.842\pm 0.004$ $0.837\pm 0.003$
Appendix G Gap Table
@@
We list the gap table in Table 12 for summarizing the current DG algorithms performance gap w.r.t FedAvg-ERM in the FL context, in particular, positive means it outperforms FedAvg-ERM, negative means it is worse than FedAvg-ERM. It can be seen that in the on the simple dataset, the best DG migrated from centralized setting is better than FedAvg-ERM. In the complete heterogeneity case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting. However, they fail in harder datasets. In the hardest setting, currently the Federated methods dealing with data heterogeneity performs the best. It is worth noting that while federated learning methods that address client heterogeneity perform better than other methods, they still fall short of achieving centralized empirical risk minimization (ERM). This highlights the need for future research and development of DG methods in the FL regime.
Table 12: Gap Table: Current Progress in solving DG in FL context
Domain Seperation
@@
Centralized adopted methods $-0.010$ $-0.047$ $+0.265$ $+0.003$ $+0.266$ $+0.003$
Federated methods $-0.040$ $+0.010$ $+0.319$ $-0.003$ $+0.266$ $-0.012$
FDG FedDG $-0.008$ $-0.010$ $+0.014$ $-0.094$
FedADG $-0.013$ $+0.060$ $+0.152$ $-0.017$
FedSR $-0.407$ $-0.104$ $-0.154$ $-0.033$ $-0.008$ $-0.124$
FedGMA $-0.225$ $-0.094$ $+0.098$ $-0.096$ $-0.145$ $-0.037$
Appendix H Training Time, Communication Rounds and local computation
@@
Table 13: Wall-clock Training time per communication (unit: s). There might be variance due to the machine’s status and resource availability.
PACS IWildCam CelebA Camelyon17 CivilComments Py150 FEMNIST
$C=100$ $C=243$ $C=100$ $C=100$ $C=100$ $C=100$ $C=100$
@@
FedAvg-ERM $143$ $6301$ $298$ $845$ $3958$ $6566$ $262$
DG Adapted IRM $147$ $6454$ $309$ $1163$ $4085$ $7089$ $297$
Fish $148$ $7072$ $312$ $1003$ $5483$ $7770$ $324$
Mixup $144$ $6294$ $302$ $933$ $264$
MMD $144$ $6663$ $312$ $963$ $4024$ $7603$ $287$
DeepCoral $144$ $6597$ $313$ $879$ $3901$ $7212$ $287$
GroupDRO $145$ $9311$ $310$ $750$ $4690$ $8121$ $307$
FL FedProx $169$ $6921$ $1219$ $4310$ $4502$ $6513$ $288$
Scaffold $167$ $6876$ $1344$ $4623$ $4421$ $6389$ $281$
FDG FedDG $352$ $32172$ $9232$ $15784$ $989$
FedADG $181$ $11094$ $6502$ $9945$ $907$
FedSR $151$ $7136$ $2267$ $6567$ $4403$ $8020$ $280$
FedGMA $143$ $6795$ $558$ $2036$ $4525$ $6545$ $261$